# Temperature Dependence of Heavy Quark Diffusion from (2+1)-flavor Lattice QCD

### v1

@@ 3.2 Normalization and Leading Order Corrections
The first normalization approach is obtained by using the leading order (LO) pQCD results for the spectral functions. We will leave the technical details to be discussed in Sec. 5, but for both, the electric and magnetic cases, the LO result is [26, 27]:
@@
where $C_{F}=(N_{c}^{2}-1)/2N_{c}$ is the Casimir factor, and $g^{2}_{{\overline{\rm MS}}}(\mu)$ is the running coupling in the ${\overline{\rm MS}}$ scheme at scale $\mu$. The transport contributions, proportional to $\kappa_{E}$ or $\kappa_{B}$, do no contribute at this order as $\kappa_{E,B}$ starts at order $g^{4}$. Then, the LO correlation function obtained by integrating Eq. (2.1) is [26]:
$G^{{\rm LO}}(\tau,T)=g^{2}_{{\overline{\rm MS}}}(\mu)C_{F}G^{\rm{norm}}(\tau,T)~{}\text{, with }~{}G^{\rm{norm}}(\tau,T)\equiv\pi^{2}T^{4}\left[\frac{\cos^{2}(\pi\tau T)}{\sin^{4}(\pi\tau T)}+\frac{1}{3\sin^{2}(\pi\tau T)}\right]\,.$ (3.7)
@@
In Refs. [51, 20, 52], the LO results for the chromo-electric and chromo-magnetic correlators have been calculated on the lattice, at finite flow times, for the improved gauge action used in this study, and for the simple discretizations of the fields given by Eqs. (3.1) and (3.2). As in our previous work [20, 52], we will use tree-level (LO) improvement to reduce the lattice discretization effects, as well as the distortions due to GF, by replacing the original lattice data $G_{E,B}^{\rm{data}}(\tau,T)$ by
@@ 3.3 Interpolations and continuum extrapolations
3.3.1 Interpolations and continuum extrapolations in the low-temperature region
For $T<205$ MeV, we have results for $m_{l}=m_{s}/20$ at three lattice spacings corresponding to $\beta=7.373,~{}7.596$ and $7.825$. Since the data for $T=153$ MeV and $\beta=7.825$ are too noisy, the lowest temperature for which we perform continuum extrapolation is $T=163$ MeV. For $\beta=7.373$ and $7.596$, we use all the temperatures shown in Tab. 1 to interpolate to $T=163$ MeV. Additionally, we perform temperature interpolations to the following values: $T=174,~{}188$ and $195$ MeV.
Then, we perform continuum extrapolations at these temperatures. Based on previous perturbative studies [52, 21, 20, 53, 19], we expect that the slopes of the extrapolations will approach zero for very large $\tau T$, with a leading term proportional to $1/(\tau T)^{2}$. Therefore, we will perform combined extrapolations for all $\tau T$, at a fixed relative flow time, using the following Ansatz:
@@ 3.3.1 Interpolations and continuum extrapolations in the low-temperature region
3.3.2 Interpolations and continuum extrapolations in the intermediate-temperature region
In the intermediate-temperature region, we have previously performed calculations at $T=220,~{}251$ and $293$ MeV with $m_{l}=m_{s}/5$. We have results in this region with $m_{l}=m_{s}/20$ with either one or two lattice spacings, as well as lattice results with $m_{l}=m_{s}/5$ for $\beta=8.4$ ($a=0.025$ fm) at two temperatures. As we have seen above, at $T=195$ MeV, the $m_{l}/20$ and $m_{s}/5$ results for the chromo-electric and chromo-magnetic correlators agree within errors. Thus, we may expect the same in the intermediate-temperature region, so it makes sense to combine both lattice results, $m_{l}=m_{s}/5$ and $m_{l}=m_{s}/20$, for the continuum extrapolations, which increases the number of lattice spacings available in each case. We find that combined continuum extrapolations work well. In addition, we can use the lattice results for $G_{E}$ and $G_{B}$ for two temperatures at $\beta=8.4$ to interpolate to $T=293$ MeV. With all these ensembles, we have five different lattice spacings at our disposal for continuum extrapolations in this temperature region. Consequently, we perform those extrapolations using Eq. (3.9), and some examples can be found in Appendix A.
3.3.3 Interpolations and continuum extrapolations in the high-temperature region
The highest temperature in the previous HotQCD lattice studies of the diffusion coefficient was $T=352$ MeV [20, 52]. However, the calculations at this temperature have been performed at a single value of the lattice spacing, $a=0.028$ fm. Therefore, in the present work, we performed calculations of $G_{E}$ and $G_{B}$ at two additional lattice spacings for $T=352$ MeV, see Tab. 2. Additionally, using the results for $\beta=8.4$, we can perform interpolations to obtain $G_{E}$ and $G_{B}$ at $T=352$ MeV with the smallest lattice spacing used in this study, $a=0.025$ fm. Therefore, altogether we have four lattice spacings for the continuum extrapolations at this temperature. Furthermore, in the present study, we extended the temperature range to higher values, considering $T=400$, $444$, $500,1000$ and $10000$ MeV. At each one of these temperatures, we performed lattice QCD calculations at three lattice spacings. Then, we used Eq. (3.9) again for the continuum extrapolations in this temperature region, and some examples can be found Appendix A.
@@ 3.4 Renormalization and flow extrapolation
The chromo-electric and chromo-magnetic correlators have to be renormalized before the spectral function can be extracted from them. The renormalization constants of these correlators, $Z_{E}$ and $Z_{B}$, are scheme and scale dependent. On the lattice, these renormalization constants in general have to be calculated before taking the continuum limit [30, 22]. Using the GF when calculating $G_{E}$ and $G_{B}$, however, automatically gives the renormalized correlation functions in the GF scheme at scale $\sqrt{8\tau_{F}}$, up to corrections proportional to powers of $\sqrt{8\tau_{F}}/\tau$. To remove these corrections (artifacts), we need to perform the $\tau_{F}\rightarrow 0$ extrapolation. For the chromo-electric correlator this extrapolation is straightforward, since at 1-loop this correlator is scheme and scale independent. This is not the case for the chromo-magnetic correlator, and the strategy of the $\tau_{F}\rightarrow 0$ extrapolation is discussed below.
@@ 3.4.1 $G_{B}$ matching factors
The chromo-magnetic correlation function is scheme and scale dependent already at 1-loop level. Since all the perturbative calculations of the chromo-magnetic spectral function used in this work are in the ${\overline{\rm MS}}$ scheme [27], it is important to know the matching between the GF and ${\overline{\rm MS}}$. This matching has been calculated at 1-loop [54, 55]:
@@
where $\gamma_{0}=3/(8\pi^{2})$ is the leading anomalous dimension and $\gamma_{\rm{E}}=0.5772156649$ is the Euler-Mascheroni constant. To improve the conversion and to account for unknown higher-loop effects and multiple logarithmic corrections, we can extend Eq. (3.10) to [52]:
@@
In this study, we convert the continuum extrapolated results for $G_{B}^{\rm{GF}}(\tau T,\mu_{F})$ to the ${\overline{\rm MS}}$ scheme using the above equations, and then perform the $\tau_{F}$ extrapolation. Unlike Ref. [52], we do not use the physical scheme for $G_{B}$ introduced in [56, 27], since comparison with weak coupling calculations and the extraction of $\kappa_{B}$ can be entirely performed within the ${\overline{\rm MS}}$ scheme.
For the numerical calculation of the matching factor, we use the $g_{{\overline{\rm MS}}}^{2}$ perturbative results up to five loops calculated with RunDec package, with $\Lambda_{\overline{\rm MS}}=0.339$ GeV and $N_{f}=3$. This way to evaluate $g_{{\overline{\rm MS}}}^{2}$ will be used throughout this paper. We need to choose the ${\overline{\rm MS}}$ renormalization scale $\mu$ such that the logarithmic corrections do not become too large. In this study, we consider four choices: $\mu=2$, $3$, $4$, and $5$ GeV, and find that for these values the logarithmic corrections are never too large to be of concern. Fig. 5 shows the matching factor $Z_{\rm{match}}(\mu,\mu_{F})$ for these four choices of $\mu$, plotted as a function of $\sqrt{8\tau_{F}}/r_{1}$. The change in $\mu$ mostly amounts to the overall shift in the value of the matching coefficient, and does not change its $\tau_{F}$-dependence in the flow time window specified in Eq. (3.5), which is indicated by the gray band in Fig. 5. As a consequence, the curvature of $G_{B}^{\rm{GF}}(\tau T,\mu_{F})$, which encodes the value of $\kappa_{B}$, remains practically unchanged by a global rescaling such that $\kappa_{B}$ turns out as $\mu$-independent within errors.
@@ 3.4.2 Flow extrapolation
Figure 7: Continuum and flow extrapolated correlators, normalized by Eq. (3.8), as a function of $\tau T$ for different temperatures. The magnetic correlator is evaluated at $\mu=4$ GeV.
Once the appropriate matching factors have been calculated, we can safely perform the flow extrapolations $\tau_{F}\rightarrow 0$. Perturbative calculation suggests that the leading flow time dependence of the chromo-electric and chromo-magnetic correlators is linear in $\tau_{F}$ [46, 47, 54]. Therefore, we will use linear form to perform the extrapolations. As discussed in Sec. 3, we will restrict the analysis within the range specified in Eq. (3.5). Within this range, the signal-to-noise ratio remains robust while the lattice artifacts, which are typically pronounced at small flow times, are effectively constrained. Additionally, it ensures a reasonable linear behavior of the extrapolation across all data sets.
For the electric correlator, the slopes of the extrapolations are expected to have a very weak dependence on $\tau T$ [47], so we choose a combined fit for all $\tau T$ at the same time. On the other hand, for the magnetic correlator, the data clearly shows that the slopes depend on $\tau T$, so we rather use independent extrapolations. In both cases we use the following approach:
@@
[…]
Finally, the continuum and zero flow extrapolated results for both correlators are shown in Fig. 7, normalized using Eq. (3.8), for the full range of temperatures considered in this work. The magnetic correlator is shown at the ${\overline{\rm MS}}$ scale $\mu=4$ GeV. The continuum and flow time extrapolated results of $G_{E}/G_{\rm norm}$ and $G_{B}/G_{\rm norm}$ have the same features as the raw lattice results, namely these ratios decrease with increasing temperature and their dependence on $\tau T$ also decreases. At the highest temperature considered, these ratios are almost $\tau$-independent. While continuum or flow extrapolations do not change the general properties of the correlators, the statistical errors increase significantly in the zero flow limit. In the left figures, we do not show the magnetic correlators for $T=163$ MeV due to the large errors, although we will still use it for the spectral reconstruction of Sec. 5.
@@ 4.1 Chromo-electric correlators and spectral functions at NLO
The chromo-electric correlator has been calculated in the literature at next-to-leading order (NLO) [57, 54, 58]. The full NLO result has no simple analytic form in general. However, it has a simple form for $\omega\ll T$ and for $\omega\gg T$ [57]. For $\omega\ll T$ we have [57]:
@@
where $\xi=1/2-\gamma_{E}+\zeta^{\prime}(2)/\zeta(2)\simeq-0.64718$ and $m_{D}^{2}\equiv g^{2}T^{2}(N_{c}+N_{f}/2)/3$ is the LO Debye mass. The leading order result for $\kappa_{E}$ was first obtained in Ref. [7]. Thus, the NLO result for $\rho_{E}$ yields the LO result for $\kappa_{E}$. For $\omega\gg T$, we have:
$\rho_{E}^{{\rm NLO}}(\omega,T)\simeq\rho_{E}^{\rm{UV},{\rm NLO}}(\omega)$ $=$ $\frac{g^{2}_{{\overline{\rm MS}}}(\mu)C_{F}}{6\pi}\omega^{3}\left\{1+\frac{g^{2}_{{\overline{\rm MS}}}(\mu)}{(4\pi)^{2}}\left[-N_{f}\left(\frac{2}{3}\ln\frac{\mu^{2}}{4\omega^{2}}+\frac{20}{9}\right)\right.\right.$
$\hskip 128.60657pt\left.\left.{}+N_{c}\left(\frac{11}{3}\ln\frac{\mu^{2}}{4\omega^{2}}+\frac{149}{9}-\frac{2\pi^{2}}{3}\right)\right]\right\}\,,$
where $g^{2}_{{\overline{\rm MS}}}(\mu)$ is the running coupling in the ${\overline{\rm MS}}$ scheme at scale $\mu$, and $N_{c}=3$, $N_{f}=3$ are the number of colors and quark flavors, respectively. In this limit, the spectral function is independent of the temperature. There was a mistake in the original calculation of $\rho_{E}$ in Ref. [57], which was corrected in Refs. [54, 59]. In the original calculation, the last term in the above equation had been reported as $-\frac{8\pi^{2}}{3}$. We also note that $\rho_{E}^{{\rm NLO}}$ is independent of $\mu$ up to higher order terms, as the explicit dependence on $\mu$ through the logarithmic correction is canceled by the $\mu$-dependence of $g^{2}_{{\overline{\rm MS}}}(\mu)$.
@@
For $N_{c}=3$, $N_{f}=3$, we obtain $\mu^{{\rm NLO}}(\omega)\simeq 0.549\omega$. Note that, due to the above mentioned error in the NLO spectral function, this scale changes drastically from its previous value of $\mu^{{\rm NLO}}(\omega)\simeq 14.7427\omega$ used in Refs. [20, 52].
The NLO spectral function interpolates smoothly between the limiting cases given by Eq. (4.1) and Eq. (4.1) [57], at least for sufficiently high temperature, where the NLO result is trustworthy, see e.g. discussion in Ref. [19]. For $\omega\sim T$, the choice $\mu=\mu^{\rm{opt}}$ may not be suitable, as the spectral function will depend on $T$. It was argued that a reasonable choice of scale in this case could be the one appearing in the effective theory of QCD at high temperatures, called EQCD [60]:
@@
While at NLO the spectral function does not depend on $\mu$, higher order corrections related to the running of the coupling are numerically important, as pointed out in Ref. [19]. Therefore, it makes sense to use the NLO result for the spectral function with the running coupling to construct the reference correlator for $G_{E}$, instead of $G_{\rm norm}$ introduced in Sec. 3.2. Therefore, we define:
@@
In Fig.8, we show the continuum and flow-time extrapolated chromo-electric correlators normalized by these LO and NLO perturbative results, $G^{{\rm LO+}}$ and $G^{{\rm NLO+}}$, respectively. For clarity, only a subset of temperatures is displayed in the figure. We note that these normalized ratios are smaller than the corresponding $G_{E}/G^{\rm norm}$ ratios shown in Fig. 7, and they exhibit a different $\tau$ dependence. On one side, the temperature dependence of the ratios at large $\tau$ is substantial, indicating that it is sensitive to the diffusion coefficient $\kappa_{E}$. In contrast, at small $\tau$, the temperature dependence is much weaker, and furthermore the ratio $G_{E}/G_{E}^{{\rm NLO+}}$ approaches one, while $G_{E}/G_{E}^{{\rm LO+}}$ remains slightly above one. This behavior demonstrates that the perturbative calculation accurately captures the behavior of $G_{E}$ at small $\tau$, as expected, and that the inclusion of NLO corrections improves upon the LO result in this region. Previous studies did not observe such an agreement with perturbative predictions [19, 20] due to the mistake in the original NLO expression for $\rho_{E}$ given in Ref. [57].
@@
At leading order, one can use the kinetic theory to obtain the following expression for $\kappa_{E}$ [7]:
@@
Unless $m_{D}\ll T$, the approximation $\Pi_{00}(p,T)\simeq m_{D}^{2}$ is a rather poor one. We can improve it by considering the leading momentum dependence of the self-energy [19]:
@@
The heavy-quark momentum diffusion coefficient was also calculated at NLO [13, 61]:

### v2

@@ 3.2 Normalization and Leading Order Corrections
The first normalization approach is obtained by using the leading order (LO) pQCD results for the spectral functions. We will leave the technical details to be discussed in Sec. 5, but for both, the electric and magnetic cases, the LO result is Caron-Huot:2009ncn ; Bouttefeux:2020ycy :
@@
where $C_{F}=(N_{c}^{2}-1)/2N_{c}$ is the Casimir factor, and $g^{2}_{{\overline{\rm MS}}}(\mu)$ is the running coupling in the $\overline{\rm MS}$} scheme at scale $\mu$. The transport contributions, proportional to $\kappa_{E}$ or $\kappa_{B}$, do no contribute at this order as $\kappa_{E,B}$ starts at order $g^{4}$. Then, the LO correlation function obtained by integrating Eq. (2.1) is Caron-Huot:2009ncn :
$G^{{\rm LO}}(\tau,T)=g^{2}_{{\overline{\rm MS}}}(\mu)C_{F}G^{\rm{norm}}(\tau,T)\,$,
with
$G^{\rm{norm}}(\tau,T)\equiv\pi^{2}T^{4}\left[\frac{\cos^{2}(\pi\tau T)}{\sin^{4}(\pi\tau T)}+\frac{1}{3\sin^{2}(\pi\tau T)}\right]\,.$ (3.7)
@@
In Refs. tuprints23185 ; Altenkort:2023oms ; Altenkort:2024spl , the LO results for the chromo-electric and chromo-magnetic correlators have been calculated on the lattice, at finite flow times, for the improved gauge action used in this study, and for the simple discretizations of the fields given by Eqs. (3.1) and (3.2). As in our previous work Altenkort:2023oms ; Altenkort:2024spl , we will use tree-level (LO) improvement to reduce the lattice discretization effects, as well as the distortions due to GF, by replacing the original lattice data $G_{E,B}^{\rm{data}}(\tau,T)$ by
@@ 3.3 Interpolations and continuum extrapolations
3.3.1 Interpolations and continuum extrapolations of correlation functions in the low-temperature region
For $T<205$ MeV, we have results for $m_{l}=m_{s}/20$ at three lattice spacings corresponding to $\beta=7.373,~7.596$ and $7.825$. Since the data for $T=153$ MeV and $\beta=7.825$ are too noisy, the lowest temperature for which we perform continuum extrapolation is $T=163$ MeV. For $\beta=7.373$ and $7.596$, we use all the temperatures shown in Tab. 1 to interpolate to $T=163$ MeV. Additionally, we perform temperature interpolations to the following values: T=174,188T=174,~188 and $195$ MeV.
Then, we perform continuum extrapolations at these temperatures. Based on previous perturbative studies Altenkort:2024spl ; Altenkort:2023eav ; Altenkort:2023oms ; Brambilla:2019oaa ; Brambilla:2020siz , we expect that the slopes of the extrapolations will approach zero for very large $\tau T$, with a leading term proportional to $1/(\tau T)^{2}$. Therefore, we will perform combined extrapolations for all $\tau T$, at a fixed relative flow time, using the following Ansatz:
@@ 3.3.1 Interpolations and continuum extrapolations of correlation functions in the low-temperature region
3.3.2 Interpolations and continuum extrapolations of correlation functions in the intermediate-temperature region
In the intermediate-temperature region, we have previously performed calculations at T=220,251T=220,~251 and $293$ MeV with $m_{l}=m_{s}/5$. We have results in this region with $m_{l}=m_{s}/20$ with either one or two lattice spacings, as well as lattice results with $m_{l}=m_{s}/5$ for $\beta=8.4$ ($a=0.025$ fm) at two temperatures. As we have seen above, at $T=195$ MeV, the $m_{l}/20$ and $m_{s}/5$ results for the chromo-electric and chromo-magnetic correlators agree within errors. Thus, we may expect the same in the intermediate-temperature region, so it makes sense to combine both lattice results, $m_{l}=m_{s}/5$ and $m_{l}=m_{s}/20$, for the continuum extrapolations, which increases the number of lattice spacings available in each case. We find that combined continuum extrapolations work well. In addition, we can use the lattice results for $G_{E}$ and $G_{B}$ for two temperatures at $\beta=8.4$ to interpolate to $T=293$ MeV. With all these ensembles, we have five different lattice spacings at our disposal for continuum extrapolations in this temperature region. Consequently, we perform those extrapolations using Eq. (3.9), and some examples can be found in Appendix A.
3.3.3 Interpolations and continuum extrapolations of correlation functions in the high-temperature region
The highest temperature in the previous HotQCD lattice studies of the diffusion coefficient was $T=352$ MeV Altenkort:2023oms ; Altenkort:2024spl . However, the calculations at this temperature have been performed at a single value of the lattice spacing, $a=0.028$ fm. Therefore, in the present work, we performed calculations of $G_{E}$ and $G_{B}$ at two additional lattice spacings for $T=352$ MeV, see Tab. 2. Additionally, using the results for $\beta=8.4$, we can perform interpolations to obtain $G_{E}$ and $G_{B}$ at $T=352$ MeV with the smallest lattice spacing used in this study, $a=0.025$ fm. Therefore, altogether we have four lattice spacings for the continuum extrapolations at this temperature. Furthermore, in the present study, we extended the temperature range to higher values, considering $T=400$, $444$, $500,1000$ and $10000$ MeV. At each one of these temperatures, we performed lattice QCD calculations at three lattice spacings. Then, we used Eq. (3.9) again for the continuum extrapolations in this temperature region, and some examples can be found in Appendix A.
@@ 3.4 Renormalization and flow extrapolation
The chromo-electric and chromo-magnetic correlators have to be renormalized before the spectral function can be extracted from them. The renormalization constants of these correlators, $Z_{E}$ and $Z_{B}$, are scheme and scale dependent. On the lattice, these renormalization constants in general have to be calculated before taking the continuum limit Christensen:2016wdo ; Banerjee:2022uge . Using the GF when calculating $G_{E}$ and $G_{B}$, however, automatically gives the renormalized correlation functions in the GF scheme at scale $\sqrt{8\tau_{F}}$, up to corrections proportional to powers of $\sqrt{8\tau_{F}}/\tau$. To remove these corrections (artifacts), we need to perform the $\tau_{F}\rightarrow 0$ extrapolation. For the chromo-electric correlator this extrapolation is straightforward, since at 1-loop this correlator is the same in the GF and $\overline{\mathrm{MS}}$ schemes and is scale independent in each of these schemes. This is not the case for the chromo-magnetic correlator, and the strategy of the $\tau_{F}\rightarrow 0$ extrapolation is discussed below.
@@ 3.4.1 $G_{B}$ matching factors
The chromo-magnetic correlation function is scheme and scale dependent already at 1-loop level. Since all the perturbative calculations of the chromo-magnetic spectral function used in this work are in the $\overline{\rm MS}$} scheme Bouttefeux:2020ycy , it is important to know the matching between the GF and $\overline{\rm MS}$}. This matching has been calculated at 1-loop delaCruz:2024cix ; Brambilla:2023vwm :
@@
where $\gamma_{0}=3/(8\pi^{2})$ is the leading anomalous dimension and $\gamma_{\rm{E}}=0.5772156649$ is the Euler-Mascheroni constant. To improve the conversion and to account for unknown higher-loop effects and multiple logarithmic corrections, we can extend Eq. (3.10) to Altenkort:2024spl :
@@
In this study, we convert the continuum extrapolated results for $G_{B}^{\rm{GF}}(\tau T,\mu_{F})$ to the $\overline{\rm MS}$} scheme using the above equations, and then perform the $\tau_{F}$ extrapolation. Unlike Ref. Altenkort:2024spl , we do not use the physical scheme for $G_{B}$ introduced in Laine:2021uzs ; Bouttefeux:2020ycy , since comparison with weak coupling calculations and the extraction of $\kappa_{B}$ can be entirely performed within the $\overline{\rm MS}$} scheme. However, an extra step is needed to obtain the final physical value $\kappa_{B}$ from these correlators. We will discuss this in more detail in Sec. 5.2.
For the numerical calculation of the matching factor, we use the $g_{{\overline{\rm MS}}}^{2}$ perturbative results up to five loops calculated with RunDec package, with $\Lambda_{\overline{\rm MS}}=0.339$ GeV and $N_{f}=3$. This way to evaluate $g_{{\overline{\rm MS}}}^{2}$ will be used throughout this paper. We need to choose the $\overline{\rm MS}$} renormalization scale $\mu$ such that the logarithmic corrections do not become too large. In this study, and for temperatures $T\neq 10$ Gev, we consider four choices: $\mu=2$, $3$, 44, and $5$ GeV, and find that for these values the logarithmic corrections are never too large to be of concern. For the highest temperature $T=10$ GeV, the logarithmic corrections are already large, so only for this temperature, we will consider a higher $\overline{\rm MS}$} scale of $\mu=50$ GeV. Fig. 5 shows the matching factor $Z_{\rm{match}}(\mu,\mu_{F})$ for these four choices of $\mu$, plotted as a function of $\sqrt{8\tau_{F}}/r_{1}$. The change in $\mu$ mostly amounts to the overall shift in the value of the matching coefficient, and does not change its $\tau_{F}$-dependence in the flow time window specified in Eq. (3.5), which is indicated by the gray band in Fig. 5. As a consequence, the curvature of $G_{B}^{\rm{GF}}(\tau T,\mu_{F})$, which encodes the value of $\kappa_{B}$, remains practically unchanged by a global rescaling such that $\kappa_{B}$ turns out as $\mu$-independent within errors.
@@ 3.4.2 Flow extrapolation
Figure 7: Continuum and flow extrapolated correlators, normalized by Eq. (3.8), as a function of $\tau T$ for different temperatures. The magnetic correlator is evaluated at $\mu=4$ GeV for $T\neq 10$ GeV, and at $\mu=50$ GeV for $T=10$ GeV.
Once the appropriate matching factors have been calculated, we can safely perform the flow extrapolations $\tau_{F}\rightarrow 0$. Perturbative calculation suggests that the leading flow time dependence of the chromo-electric and chromo-magnetic correlators is linear in $\tau_{F}$ Eller:2018yje ; Eller:2021qpp ; delaCruz:2024cix . Therefore, we will use linear form to perform the extrapolations. As discussed in Sec. 3, we will restrict the analysis within the range specified in Eq. (3.5). Within this range, the signal-to-noise ratio remains robust while the lattice artifacts, which are typically pronounced at small flow times, are effectively constrained. Additionally, it ensures a reasonable linear behavior of the extrapolation across all data sets.
For the electric correlator, the slopes of the extrapolations are expected to have a very weak dependence on $\tau T$ Eller:2021qpp , so we choose a combined fit for all $\tau T$ at the same time. On the other hand, for the magnetic correlator, the data clearly shows that the slopes depend on $\tau T$, so we rather use independent extrapolations. In both cases we use the following approach:
@@
Fig. 6 shows the flow extrapolation of $G_{E,B}$ for two different temperature values: $T=195$ and $400$ MeV. For the magnetic correlator, we show the extrapolation at the $\overline{\rm MS}$} scale $\mu=4$ GeV. For most temperatures, the choice of scale acts only as a multiplicative factor, and the slopes of the extrapolation remain quite independent of $\mu$. As discussed in Sec. 3.2, this will lead to a very weak dependence of the diffusion coefficient on the choice of the renormalization scale, since $\kappa_{B}$ is mainly determined by the curvature of the correlators. In both, the electric and the magnetic cases, the extrapolations give reasonable $\chi^{2}/\text{d.o.f}$ over the whole temperature range we consider.
Finally, the continuum and zero flow extrapolated results for both correlators are shown in Fig. 7, normalized using Eq. (3.8), for the full range of temperatures considered in this work. The magnetic correlator is shown at the $\overline{\rm MS}$} scale $\mu=4$ GeV. The continuum and flow time extrapolated results of $G_{E}/G_{\rm norm}$ and $G_{B}/G_{\rm norm}$ have the same features as the raw lattice results, namely these ratios decrease with increasing temperature and their dependence on $\tau T$ also decreases. At the highest temperature considered, these ratios are almost $\tau$-independent. While continuum or flow extrapolations do not change the general properties of the correlators, the statistical errors increase significantly in the zero flow limit. In the right panel, we do not show the magnetic correlators for $T=163$ MeV due to the large errors, although we will still use it for the spectral reconstruction of Sec. 5.
@@ 4.1 Chromo-electric correlators and spectral functions at NLO
The chromo-electric correlator has been calculated in the literature at next-to-leading order (NLO) Burnier:2010rp ; delaCruz:2024cix ; Scheihing-Hitschfeld:2023tuz . The full NLO result has no simple analytic form in general. However, it has a simple form for $\omega\ll T$ and for $\omega\gg T$ Burnier:2010rp . For $\omega\ll T$ we have Burnier:2010rp :
@@
where $\xi=1/2-\gamma_{E}+\zeta^{\prime}(2)/\zeta(2)\simeq-0.64718$ and $m_{D}^{2}\equiv g^{2}T^{2}(N_{c}+N_{f}/2)/3$ is the LO Debye mass. The leading order result for $\kappa_{E}$ was first obtained in Ref. Moore:2004tg . Thus, the NLO result for $\rho_{E}$ yields the LO result for $\kappa_{E}$. For $\omega\gg T$, we have:
[…]
$\left.\left.{}+N_{c}\left(\frac{11}{3}\ln\frac{\mu^{2}}{4\omega^{2}}+\frac{149}{9}-\frac{2\pi^{2}}{3}\right)\right]\right\}\,$,
where $g^{2}_{{\overline{\rm MS}}}(\mu)$ is the running coupling in the $\overline{\rm MS}$} scheme at scale $\mu$, and $N_{c}=3$, $N_{f}=3$ are the number of colors and quark flavors, respectively. In this limit, the spectral function is independent of the temperature. There was a mistake in the original calculation of $\rho_{E}$ in Ref. Burnier:2010rp , which was corrected in Refs. delaCruz:2024cix ; Sambataro:2023tlv . In the original calculation, the last term in the above equation had been reported as $\frac{8\pi^{2}}{3}$. We also note that $\rho_{E}^{{\rm NLO}}$ is independent of $\mu$ up to higher order terms, as the explicit dependence on $\mu$ through the logarithmic correction is canceled by the $\mu$-dependence of $g^{2}_{{\overline{\rm MS}}}(\mu)$.
@@
For $N_{c}=3$, $N_{f}=3$, we obtain $\mu^{{\rm NLO}}(\omega)\simeq 0.549\omega$. Note that, due to the above-mentioned error in the NLO spectral function, this scale changes drastically from its previous value of $\mu^{{\rm NLO}}(\omega)\simeq 14.7427\omega$ used in Refs. Altenkort:2023oms ; Altenkort:2024spl .
The NLO spectral function interpolates smoothly between the limiting cases given by Eq. (4.1) and Eq. (4.1) Burnier:2010rp , at least for sufficiently high temperature, where the NLO result is trustworthy, see e.g. discussion in Ref. Brambilla:2020siz . For $\omega\sim T$, the choice $\mu=\mu^{\rm{opt}}$ may not be suitable, as the spectral function will depend on $T$. It was argued that a reasonable choice of scale in this case could be the one appearing in the effective theory of QCD at high temperatures, called EQCD Kajantie:1997tt :
@@
While at NLO the spectral function does not depend on $\mu$, higher order corrections related to the running of the coupling are numerically important, as pointed out in Ref. Brambilla:2020siz . Therefore, it makes sense to use the NLO result for the spectral function with the running coupling to construct the reference correlator for $G_{E}$, instead of $G_{\rm norm}$ introduced in Sec. 3.2. Therefore, we define:
@@
In Fig.8, we show the continuum and flow-time extrapolated chromo-electric correlators normalized by these LO and NLO perturbative results, $G^{{\rm LO+}}$ and $G^{{\rm NLO+}}$, respectively. For clarity, only a subset of temperatures is displayed in the figure. We note that these normalized ratios are smaller than the corresponding $G_{E}/G^{\rm norm}$ ratios shown in Fig. 7, and they exhibit a different $\tau$ dependence. On one side, the temperature dependence of the ratios at large $\tau$ is substantial, indicating that it is sensitive to the diffusion coefficient $\kappa_{E}$. In contrast, at small $\tau$, the temperature dependence is much weaker, and furthermore the ratio $G_{E}/G_{E}^{{\rm NLO+}}$ approaches one, while $G_{E}/G_{E}^{{\rm LO+}}$ remains slightly above one. This behavior demonstrates that the perturbative calculation accurately captures the behavior of $G_{E}$ at small $\tau$, as expected, and that the inclusion of NLO corrections improves upon the LO result in this region. Previous studies did not observe such an agreement with perturbative predictions Brambilla:2020siz ; Altenkort:2023oms due to the mistake in the original NLO expression for $\rho_{E}$ given in Ref. Burnier:2010rp .
@@
At leading order, one can use the kinetic theory to obtain the following expression for $\kappa_{E}$ Moore:2004tg :
@@
Unless $m_{D}\ll T$, the approximation $\Pi_{00}(p,T)\simeq m_{D}^{2}$ is a rather poor one. We can improve it by considering the leading momentum dependence of the self-energy Brambilla:2020siz :
@@
The heavy-quark momentum diffusion coefficient was also calculated at NLO Caron-Huot:2007rwy ; Caron-Huot:2008dyw :